# Identifying Best Candidates for Busbar Splitting

### v1

@@ Abstract
Rising electricity demand and the growing integration of renewables are intensifying congestion in transmission grids. Grid topology optimization through busbar splitting (BuS) and optimal transmission switching can alleviate grid congestion and reduce the generation costs in a power system. However, BuS optimization requires a large number of binary variables, and analyzing all the substations for potential new topological actions is computationally intractable, particularly in large grids. To tackle this issue, we propose a set of metrics to identify and rank promising candidates for BuS, focusing on finding buses where topology optimization can reduce generation costs. To assess the effect of BuS on the identified buses, we use a combined mixed-integer convex-quadratic BuS model to compute the optimal topology and test it with the non-linear non-convex AC optimal power flow (OPF) simulation to show its AC feasibility. By testing and validating the proposed metrics on test cases of different sizes, we show that they are able to identify busbars that reduce the total generation costs when their topology is optimized. Thus, the metrics enable effective selection of busbars for BuS, with no need to test every busbar in the grid, one at a time.
Keywords: Busbar Splitting, Grid Topology Optimization, Metrics, Topological Actions, Transmission Grids.
1 Introduction and motivation
@@
With the increasing urgency to decrease global CO2 emissions, societies are becoming increasingly electrified. As a result, electricity demand is projected to increase significantly [1]. However, due to the limited transmission capacity and the distributed nature of newly-installed renewable energy sources (RES) capacity [2], power systems worldwide are facing increasing congestion [1]. In both Europe [3] and the US [4], congestion costs have seen a substantial increase in the last years, reaching e.g. 2.774 bn€ in 2024 in Germany [5], and 8-13.5 bn\$ in 2022 and 2023 for all the US Independent System Operators [4]. Currently, power grids are considered an important bottleneck for a carbon-free society, as the deployment of RES requires additional transmission capacity [1] and significant RES curtailment will take place without effective grid expansion projects [6]. At the same time, expanding the grid requires worldwide investments in the range of hundreds of billions €/\$ per year [7]. Besides economic reasons, grid expansion projects are often delayed due to technical (supply chain issues, lack of workforce, lack of manufacturing capacity, bottlenecks with critical elements, etc. [8]) and non-technical (public opposition [9]) issues. Because of all these reasons, system operators aim to maximize the transmission capacity in the existing grid and reduce congestion. Congestion management actions such as power flow control can be performed with phase-shifting transformers and high-voltage direct current links. A broader term to describe these technologies is “grid-enhancing technologies” [10], which also includes flexible AC transmission technologies and dynamic line rating. Although the attention surrounding these technologies is increasing [11], in Europe, they are still primarily applied to maximize cross-border flows, while additional costly remedial measures are employed to address internal grid congestion.
@@
Consequently, several initiatives in Europe [12] and the US [11, 13] aim to improve the utilization of existing grid infrastructure by actively optimizing and reconfiguring the network topology to mitigate congestion and maximize transfer capacity. However, the available topological actions, such as OTS and substation reconfiguration with BuS are often restricted to well-established, expert-driven measures that system operators have employed for decades, and may not take into account the distributed or variable nature of recently installed RES capacity. As a result, potentially effective topological actions remain unidentified, and systematic methodologies to discover them are yet to be established in the scientific literature. In particular, while the most common topological actions in industry include substation reconfiguration through BuS, academic papers investigating topology optimization with BuS are relatively new, in part because Bus is computationally highly challenging. Every busbar considered in the topology optimization problem introduces a new degree of freedom and thus has the potential to reduce generation costs compared to a standard OPF, but representing multiple possible splitting configurations requires the use of binary variables and thus adds significant computational complexity. When implementing topology optimization for larger grids, it is therefore necessary to pre-screen busbars to identify promising candidate busbars for BuS. This paper addresses this issue by proposing metrics to identify efficient topological actions in high voltage grids based on the results of optimal power flow (OPF) simulations. Although several metrics (discussed later) have been proposed in the literature, no prior work has systematically assessed their effectiveness in identifying the most suitable busbars for applying topological actions. In a rapidly evolving power system, such metrics can support system operators in identifying new impactful topological actions, ultimately maximizing the grid’s power transfer capacity.
1.1 Related work and contributions
@@
The concept of topology optimization using optimal transmission switching (OTS) and busbar splitting (BuS) has been utilized for decades to manage contingencies [14], reduce the total generation costs (OTS [15, 16] and BuS [17]), relieve grid congestion, and plan maintenance. While there is a considerable body of literature on OTS [18], its extensions [19] and application related to contingency analysis [16], the use of BuS optimization to achieve reductions in the total generation costs is fairly recent [20, 21, 22, 23] and has been applied to hybrid AC/DC grids, too [24, 25, 26]. Furthermore, the first ideas to filter and select the most relevant network elements to be switched were developed decades ago for [27, 28]. Most literature is based on the concept of locational marginal price (LMP), defined as “the marginal cost of delivering one more unit of electricity to a specific bus, taking into account transmission losses and network constraints” [29]. LMPs depend on the location where electricity is generated or consumed, and are usually higher in areas that normally import power and lower in areas that export power [30]. From a mathematical perspective, LMPs can be derived as the Lagrange multiplier associated with the nodal balance constraint in the OPF formulation [30].
@@
References  [31] and [32] propose a heuristic to select the most relevant lines to be switched in the OTS problem based on the difference in LMPs between the two ends of a branch, multiplied by the power flow through the branch. Additional sensitivity criteria for line selection are investigated in [33]. Moreover, potential remedial actions are identified by studying congestion zones in [34] for the OTS problem and in [20, 35] for the BuS problem. [22] extends the methodology in [20] by adding a pre-screening step that uses LMPs associated with active and reactive power and only considers substations with four or more connected lines to identify relevant busbars for OTS. Other references [36, 37] develop heuristics for selecting the most relevant busbars in BuS based on the load margin [36] and for selecting the best breakers in node-breaker modelling [37] using i) the maximum increase in the congestion rent, ii) the reduction in total generation costs and iii) the largest absolute value of LMP difference across any line in the system. The latter metric is similar to [31], but without considering the power through the lines. Finally,  [23] selects buses with ten or more connected components as candidates for topology optimization. In previous work [24, 26], we observed that only certain busbars lead to a reduction in generation costs when only one busbar was allowed to be split at a time. However, if applied to a large system, having to exhaustively checking all busbars adds significant computational cost. The goal in this paper is therefore to identify the small subset of buses that will reduce generation cost if we optimize their topology.
@@
Table 1 summarizes the most relevant metrics emerging from the literature, and whether they were related to OTS or BuS in the aforementioned publications. The new metrics proposed in this work are introduced later in Section 3.
@@
Table 1: Overview of common metrics in the literature related to optimal transmission switching (OTS) and busbar splitting (BuS). “LMP” stands for “locational marginal price”. $\Delta_{LMP}$” stands for the difference in LMPs over a branch.
@@
Congestion zones identification [34, 35] [20, 22]
@@
1.2 Contributions
@@
  • Identifying and validating metrics to select relevant busbar candidates to perform busbar splitting on AC grids ranging from 39 to thousands of buses, with the final objective of reducing the total generation costs. The pre-emptive selection of relevant candidates for busbar splitting considerably reduces the computational time compared to splitting one bus at a time.
@@
The remainder of this work is organized as follows. Section 2 describes the existing topology optimization model used in the the paper, while Section 3 describes the proposed metrics for selecting the best candidates for BuS. Section 4 discusses the results, with Section 5 concluding the paper, and highlighting possible avenues for future work.
2 Preliminaries: Topology Optimization with Busbar Splitting
@@
This paper leverages the combined BuS model introduced in [24, 26] to solve the topology optimization problem with an LPAC-OPF formulation. This model gives rise to a mixed-integer, convex-quadratic optimization problem that can be solved relatively efficiently compared with a formulation that uses the full non-convex, non-linear AC power flow constraints, while also providing results that are typically AC-feasible [24, 26]. Once we obtain the optimized topology, we evaluate the cost reduction by solving the AC OPF from [38] with the original and optimized topologies. For brevity, we refer the reader to  [24, 26, 38] for a thorough explanation of the BuS model and the detailed AC-OPF formulation. Here, we focus on explaining the logical constraints used to represent the BuS actions in the model to highlight their inherent complexity.
Fig. 1 provides an example of what is needed to represent a bus with two branches, a generator and a load.
Figure 1: Steps behind the topology optimization model proposed in [24, 26]. The selected busbar $i$ (left) is split into two buses $i$ and $i^{\prime}$, connected through a busbar coupler. The open/close position of this busbar coupler is represented by the binary variable $ZIL_{ii^{\prime}}$. Each network element that was connected to busbar $i$ in the input topology is linked to an auxiliary bus (named $m,n,o$ and $p$ in the figure) and can be connected to either one (ii) or the other part ($i^{\prime}$) of the split busbar through a switch (center). After the optimization, the inactive switches are removed to yield the new topology (right).
@@
To represent BuS actions for busbar $i$ in Fig. 1 (left), we remove all grid elements connected to the busbar and instead connect them to new auxiliary busbars. Then, the original bus is split into two buses $i$ and $i^{\prime}$. For each component, we introduce a pair of switches $sw_{\upsilon mi}^{ac}$} and $sw_{\kappa mi^{\prime}}^{ac}$} to represent which part of the split busbar the component can be connected to. Each component cannot be connected to more than one bus, but can be completely disconnected. We also introduce a binary variable $sw_{ZILii^{\prime}}^{ac}$} to represent the open/close status of the busbar coupler, which connects the busbars $i$ and $i^{\prime}$. If the busbar coupler $sw_{ZILii^{\prime}}^{ac}$} is closed, the two parts of the AC busbar are connected, and AC buses $i$ and $i^{\prime}$ have the same voltage values. This reasoning is formulated with constraints (M1.1) to (M1.7) in Model 1, applied to a set of switches $\mathcal{T}^{\text{sw,ac}}$ and busbar couplers $\mathcal{T}^{\text{ZIL,ac}}$, with each switching element represented by a binary variable (M1.1) [^1] The nomenclature used for the constraints is based on PowerModels.jl [38] (OPF) and PowerModelsTopologicalActions.jl [24] (OTS/BuS)..
Model 1: Topology optimization constraints
$z^{sw,ac}_{\upsilon mi}\in\{0,1\}\quad\forall\upsilon mi\in\mathcal{T}^{\text{sw,ac}}\cup\mathcal{T}^{\text{ZIL,ac}}$ (M1.1)
$(1-z^{sw,ac}_{\upsilon mi})\cdot M_{\theta}\leq\theta_{m}-\theta_{i}\leq(1-z^{sw,ac}_{\upsilon mi})\cdot M_{\theta}$ (M1.2)
$\forall\upsilon mi\in\mathcal{T}^{\text{sw,ac}}\cup\mathcal{T}^{\text{ZIL,ac}}$
$(1-z^{sw,ac}_{\upsilon mi})\cdot M_{m}\leq U^{m}_{m}-U^{m}_{i}\leq(1-z^{sw,ac}_{\upsilon mi})\cdot M_{m}$ (M1.3)
$\forall\upsilon mi\in\mathcal{T}^{\text{sw,ac}}\cup\mathcal{T}^{\text{ZIL,ac}}$
$z^{sw,ac}_{\upsilon mi}\cdot\underline{P}^{sw,ac}_{\upsilon mi}\leq P^{sw,ac}_{\upsilon mi}\leq z^{sw,ac}_{\upsilon mi}\cdot\overline{P}^{sw,ac}_{\upsilon mi}$ (M1.4)
$\forall\upsilon mi\in\mathcal{T}^{\text{sw,ac}}\cup\mathcal{T}^{\text{ZIL,ac}}$
$z^{sw,ac}_{\upsilon mi}\cdot\underline{Q}^{sw,ac}_{\upsilon mi}\leq Q^{sw,ac}_{\upsilon mi}\leq z^{sw,ac}_{\upsilon mi}\cdot\overline{Q}^{sw,ac}_{\upsilon mi}$ (M1.5)
$\forall\upsilon mi\in\mathcal{T}^{\text{sw,ac}}\cup\mathcal{T}^{\text{ZIL,ac}}$
@@
$z^{sw,ac}_{\upsilon mi}+z^{sw,ac}_{\kappa mi^{\prime}}\leq 1$$\forall(\upsilon mi,\kappa mi^{\prime})\in\mathcal{T}^{\text{sw,ac}}$ (M1.6)
$z^{sw,ac}_{\kappa mi^{\prime}}\leq(1-z^{sw,ac}_{ZILii^{\prime}})$$\forall(\kappa mi,ZILii^{\prime})\in\mathcal{T}^{\text{sw,ac}}\cup\mathcal{T}^{\text{ZIL,ac}}$ (M1.7)

### v2

@@ Abstract
Rising electricity demand and the growing integration of renewables are intensifying congestion in transmission grids. Grid topology optimization through busbar splitting (BuS) and optimal transmission switching can alleviate grid congestion and reduce the generation costs in a power system. However, BuS optimization requires a large number of binary variables, and analyzing all the substations for potential new topological actions is computationally intractable, particularly in large grids. To tackle this issue, we propose a set of metrics to identify and rank promising candidates for BuS, focusing on finding buses where topology optimization can reduce generation costs. To assess the effect of BuS on the identified buses, we use a combined mixed-integer convex-quadratic BuS model to compute the optimal topology and test it with the non-linear non-convex AC optimal power flow (OPF) simulation to show its AC feasibility and generation cost reduction compared to the AC-OPF simulations. By testing and validating the proposed metrics on test cases of different sizes, we show that they are able to identify busbars that reduce the total generation costs when their topology is optimized. Thus, the metrics enable effective selection of busbars for BuS, with no need to test every busbar in the grid, one at a time.
I Introduction and motivation
@@
With the increasing urgency to decrease global CO2 emissions, societies are becoming increasingly electrified. As a result, electricity demand is projected to increase significantly [23]. However, due to the limited transmission capacity and the distributed nature of newly-installed renewable energy sources (RES) capacity [26], power systems worldwide are facing increasing congestion [23]. In both Europe [1] and the US [39, 38], congestion costs have seen a substantial increase in recent years, reaching e.g. 2.774 bn€ in 2024 in Germany [40], and 13.5-8.33 bn\$ in 2023 and 2024 for all the US Independent System Operators [39, 38]. Currently, power grids are considered an important bottleneck for a carbon-free society, and the deployment of RES requires additional transmission capacity [23]. Significant RES curtailment will take place without effective grid expansion projects aimed to increase the grid transmission capacity [13]. At the same time, expanding the grid requires worldwide investments in the range of hundreds of billions €/\$ per year [24]. Besides economic reasons, grid expansion projects are often delayed due to technical (supply chain issues, lack of workforce, lack of manufacturing capacity, bottlenecks with critical elements, etc. [25]) and non-technical (public opposition [35]) issues. Because of all these reasons, system operators are currently aiming to maximize the transmission capacity in the existing grid to reduce existing congestion. For example, congestion management actions such as power flow control can be performed, among the others, with phase-shifting transformers and high-voltage direct current links. A broader term to describe these technologies is “grid-enhancing technologies” [31], which also includes flexible AC transmission technologies and dynamic line rating. Although the attention surrounding these technologies is increasing [12], in Europe, they are still primarily applied to maximize cross-border flows, while additional costly remedial measures are employed to address internal grid congestion.
@@
Consequently, several initiatives in Europe [7] and the US [12, 43] aim to improve the utilization of existing grid infrastructure by actively optimizing and reconfiguring the network topology to mitigate congestion and maximize transfer capacity. However, the available topological actions, such as OTS and substation reconfiguration with BuS are often restricted to well-established, expert-driven measures that system operators have employed for decades, even though the power flows through transmission grids are rapidly changing. As a result, potentially effective topological actions remain unidentified, and systematic methodologies to discover them are yet to be established in the scientific literature. In particular, while the most common topological actions in industry include substation reconfiguration through BuS, academic papers investigating topology optimization with BuS are relatively new, in part because BuS is computationally highly challenging. Every busbar considered in the topology optimization problem introduces a new degree of freedom and thus has the potential to reduce generation costs compared to a standard OPF by solving congestion, therefore allowing the cheaper generators to be used to their maximum rated power, but representing multiple possible splitting configurations requires the use of binary variables and thus adds significant computational complexity. When implementing topology optimization for larger grids, it is therefore necessary to pre-screen busbars to identify promising candidate busbars for BuS. This paper addresses this issue by proposing metrics to identify efficient topological actions in transmission grids based on the results of optimal power flow (OPF) simulations. Although several metrics (discussed later) have been proposed in the literature, no prior work has systematically assessed their effectiveness in identifying the most suitable busbars for applying topological actions through BuS. In a rapidly evolving power system, such metrics can support system operators in identifying new impactful topological actions, ultimately maximizing the grid’s power transfer capacity.
I-A Related work and contributions
@@
The concept of topology optimization using optimal transmission switching (OTS) and busbar splitting (BuS) has been utilized for decades to manage contingencies [4], reduce the total generation costs (OTS [14, 17] and BuS [20]), relieve grid congestion, and plan maintenance. While the first ideas to filter and select the most relevant network elements to be switched were developed decades ago for [30, 29], and there is a considerable body of literature on OTS [18], its extensions [16] and application related to contingency analysis [17], the proposal to use BuS to optimize the substation configuration and achieve reductions in the total generation costs is fairly recent [21, 22, 19, 34] and has been applied to hybrid AC/DC grids, too [6, 33, 5, 41].
Most literature is based on the concept of locational marginal price (LMP), defined as “the marginal cost of delivering one more unit of electricity to a specific bus, taking into account transmission losses and network constraints” [32]. LMPs depend on the location where electricity is generated or consumed, and are usually higher in areas that normally import power and lower in areas that export power [28]. From a mathematical perspective, LMPs can be derived as the Lagrange multiplier associated with the nodal balance constraint in the OPF formulation [28].
@@
References  [36] and [15] propose a heuristic to select the most relevant lines to be switched in the OTS problem based on the difference in LMPs between the two ends of a branch, multiplied by the power flow through the branch. Additional sensitivity criteria for line selection are investigated in [37]. Moreover, potential remedial actions are identified by studying congestion zones in [27] for the OTS problem and in [21, 44] for the BuS problem. [19] extends the methodology in [21] by adding a pre-screening step that uses LMPs associated with active and reactive power and only considers substations with four or more connected lines to identify relevant busbars for OTS. Other references [42, 3] develop heuristics for selecting the most relevant busbars in BuS based on the load margin [42] and for selecting the best breakers in node-breaker modelling [3] using i) the maximum increase in the congestion rent, ii) the reduction in total generation costs and iii) the largest absolute value of LMP difference across any line in the system. The latter metric is similar to [36], but without considering the power through the lines. Finally,  [34] selects buses with ten or more connected components as candidates for topology optimization. The existing above-mentioned metrics to select the most relevant elements for grid topology optimization mainly rely on single $\Delta_{LMP}$s through a branch connected to a given busbar, without including $\Delta_{LMP}$ patterns throughout all the branches connected to it. In addition, except for [19], they rely on the linear DC approximation of the OPF problem, without considering the grid’s reactive power and voltage magnitude levels. These factors are becoming increasingly important in modern converter-dominated power systems, and their mismanagement could lead to events such as the Iberian blackout in April 2025 [11].
In previous work [6, 5], we observed that only certain busbars lead to a reduction in generation costs when only one busbar was allowed to be split at a time. However, if applied to a large system, having to exhaustively check all busbars adds a significant computational cost to the optimization problem. The goal in this paper is therefore to identify a small subset of busbars that will reduce generation cost if we optimize their topology. These busbars are selected a priori through the proposed metrics, based on the grid topology and on the results of the AC-OPF formulation.
@@
Table I summarizes the most relevant metrics emerging from the literature, and whether they were related to OTS or BuS in the aforementioned publications. The new metrics proposed in this work are introduced later in Section III.
@@
TABLE I: Overview of common metrics in the literature related to optimal transmission switching (OTS) and busbar splitting (BuS). “LMP” stands for “locational marginal price”. $\Delta_{LMP}$” stands for the difference in LMPs over a branch.
@@
Congested lines/zones identification [27] [21, 19, 44]
@@
Figure 1: Steps behind the topology optimization model proposed in [6, 5]. The selected busbar $i$ (left) is split into two buses $i$ and $i^{\prime}$, connected through a busbar coupler $ZIL_{ii^{\prime}}$. The open/close position of this busbar coupler $ZIL_{ii^{\prime}}$ is represented in the optimization model by a binary variable. Each network element that was connected to busbar $i$ in the input topology is linked to an auxiliary bus (named $m,n,o$ and $p$ in the figure) and can be connected to either one (ii) or the other part ($i^{\prime}$) of the split busbar through a switch (center). Each switch is also represented in the model through a binary variable. After the optimization, the inactive switches are removed to yield the new topology (right).
I-B Contributions
@@
  • Identifying and validating metrics to select relevant busbar candidates to perform busbar splitting on AC grids ranging from 39 to thousands of buses, with the final objective of reducing the total generation costs. The pre-emptive selection of relevant candidates for busbar splitting identifies zones where grid topology optimization is relevant without the need to investigate the whole grid.
@@
The remainder of this work is organized as follows.Section II describes the existing topology optimization model used in the paper, while Section III describes the proposed metrics for selecting the best candidates for busbar splitting. Section IV discusses the results, withSection V concluding the paper, and highlighting possible avenues for future work.
II Preliminaries: Topology Optimization with Busbar Splitting
@@
This paper is based on the grid topology optimization model introduced in [6, 5], which includes BuS of selected busbars based on an LPAC-OPF formulation. This model gives rise to a mixed-integer, convex-quadratic optimization problem that can be solved efficiently compared with a formulation that uses the full non-convex, non-linear AC power flow constraints, while also providing results that are typically AC-feasible [6, 5].
Once we obtain the optimized topology, we evaluate the cost reduction by solving the AC-OPF from  [8] with the original and optimized topologies. For brevity, we refer the reader to  [8] for a thorough explanation of the detailed LPAC- and AC-OPF formulations, while here we add the LPAC-BuS model from [6], and we focus on explaining the logical constraints used to represent the BuS actions in the model to highlight their inherent complexity.
II-A Intuition behind busbar splitting
@@
Fig. 1 provides an example of a busbar with two branches, a generator and a load. To represent BuS actions for busbar $i$ in Fig. 1 (left), we remove all grid elements connected to the busbar and instead connect them to new auxiliary busbars. Then, the original bus is split into two buses $i$ and $i^{\prime}$. For each component, we introduce a pair of switches $sw_{\upsilon mi}^{ac}$} and $sw_{\kappa mi^{\prime}}^{ac}$} to represent which part of the split busbar the component can be connected to. Each component cannot be connected to more than one bus, but can be completely disconnected. We also introduce a binary variable $sw_{ZILii^{\prime}}^{ac}$} to represent the open/close status of the busbar coupler, which connects the busbars $i$ and $i^{\prime}$. If the busbar coupler $sw_{ZILii^{\prime}}^{ac}$} is closed, the two parts of the AC busbar are connected, and AC buses $i$ and $i^{\prime}$ have the same voltage angle and magnitude values. This reasoning is translated into the optimization model from [6, 5] in the following Section.
II-B Grid topology optimization model
The nomenclature of the network components, topologies and connectivities used throughout the Section are included in Table II. Note that by using ‘reverse’ topologies for the branches, power flows in both directions can be represented in the OPF and BuS models [8].
TABLE II: Network components, topologies and connectivities and nomenclature of each network component included in the optimization model. “BuS” stands for “busbar splitting”.
Network components
$\mathcal{I}$ Set of nodes
$\mathcal{L}$ Set of branches
$\mathcal{C}$ Set of busbar couplers
$\mathcal{G}$ Set of generators
$\mathcal{M}$ Set of loads
$\mathcal{I^{\prime}}$ Set of auxiliary nodes for BuS
$\mathcal{S}$ Set of auxiliary switches for BuS
Topologies and connectivities
$\mathcal{T}^{ac}$ Branch topologies
$\mathcal{T}^{ac,rev}$ Reverse branch topologies
$\mathcal{T}^{\text{sw,ac}}$ Switch topologies
$\mathcal{T}^{\text{ZIL,ac}}$ Busbar coupler topologies
$\mathcal{T}^{\text{load}}$ Load connectivity
$\mathcal{T}^{\text{gen}}$ Generator connectivity
Nomenclature of each network component
$lij~\in~\mathcal{T}^{ac}$$\subseteq~\mathcal{L}$$\times~\mathcal{I}$$\times~\mathcal{I}$ Branches
@@ II-B Grid topology optimization model
$\upsilon ii^{\prime}~\in~\mathcal{T}^{\text{ZIL,ac}}\subseteq~\mathcal{C}$$\times~\mathcal{I}$$\times~\mathcal{I^{\prime}}$ Busbar couplers
$\upsilon mi~\in~\mathcal{T}^{\text{sw,ac}}\subseteq~\mathcal{S}$$\times~\mathcal{I}$$\times~\mathcal{I^{\prime}}$ Auxiliary switches for BuS
$lji~\in~\mathcal{T}^{ac,rev}$$\subseteq~\mathcal{L}$$\times~\mathcal{I}$$\times~\mathcal{I}$ Reverse Branches
$mi~\in~\mathcal{T}^{\text{load}}$ Loads
$gi~\in~\mathcal{T}^{\text{gen}}$ Generators
The LPAC-BuS model is represented by the equations (M1.1)-(M1.18), as shown in Model 1. The objective function in (M1.1) consists of i) minimizing the total costs of the $\mathcal{G}$ set of generators, where $P_{k}^{g}$ is the active power injection, and ii) adding a (small) penalty term $c_{ZIL}$ to the busbar couplers $\mathcal{C}$ to split a busbar only if it actually brings economic benefit. All the generators $k$ in the network are assigned cost parameters $c_{2k}$ ($W^{2}$), $c_{1k}$ (\$/W), $c_{0k}$ (\$). Note that the quadratic cost term of each generator is assigned to zero.
Eq. (M1.2) sets the voltage angle of the reference bus to zero while (M1.3) and (M1.4) limit respectively the voltage magnitude and angles of the other buses between minimum ($\underline{U}^{m}_{i},\underline{\theta}_{i}$) and maximum ($\overline{U}^{m}_{i},\overline{\theta}_{i}$) values. Similarly, the generator setpoints are bounded by $\underline{S}^{g}_{k}$ and $\overline{S}^{g}_{k}$ in (M1.6). The power balance for each node (or substation) is included in (M1.5), with $S^{m}_{l}$ being the nodal demand. The term $Y_{i}^{s}\cdot(1+2\phi_{i})$ refers to the power absorbed by shunt elements connected to the node $i$, which are considered negligible in this paper. $\phi_{i}$ indicates the voltage magnitude change in the LPAC formulation [10]. The active and reactive power flow through the branches are modeled through (M1.7) and (M1.8), with each branch being represented by a $\pi$-section model [8]. These active and reactive powers are constrained by (M1.9) and (M1.10), while the piecewise linear approximation of the cosine term $cs$ of $cos(\theta_{n}-\theta_{m})$ for the LPAC formulation [10] is expressed in (M1.11).